The core elements of the Growth Employment and Redistribution (GEAR) strategy of the South African government in 1996, under the leadership of the then finance minister Trevor Manuel were amongst other things:


 budget reform to strengthen the redistributive thrust of expenditure
 monetary policy to prevent a resurgence of inflation
 a reduction in tariffs to contain input prices and facilitate industrial restructuring, compensating partially for the exchange rate depreciation


1.1 With reference to the above, identify the macroeconomic objectives in these elements.
1.2 Identify and define the macroeconomic variables that can be used to measure whether the strategy was successful or not.
1.3 With each tool in 1.2, provide a detailed explanation on how it can be measured
Answer #1

1.1

The Macroeconomic Objectives are:

  • Redistributing Government Expenditure
  • Using the Budget Fund Efficiently
  • Reducing Trade Deficit
  • Control Inflation
  • Stabilising GDP
  • Minimising Inequality
  • Stable Exchange Rate

1.2

To check whether the strategy was successful or not the following macroeconomic variables can be evaluated :

GDP growth : We can see the growth of GDP over the previous term(quarter/year) , a positive growth represents a growing economy .

Inflation Rate : A increasing inflation rate signifies that the prices of goods and services are increasing within a country and a high inflation rate is undesirable , a low inflation rate is what a growing economy should have .

Exchange Rate : A strong currency value in the currency exchange market represents a growth i.e strengthened trade position .

BoP : Balance of Payments indicates the net current account(net exports-net imports) and the net capital+financial account and gives a glimpse into the trade position of the country/economy .

GDP per capita : It is used as an indicator to see how much on average an individual earns .It is given by : GDP/population

Higher the GDP per capita more is the income per individual i.e more spending power .

1.3

The GDP growth is measured by the following formula :
GDP(new term)-GDP(old term)/GDP(old term)

For example we need to calculate GDP growth for year 2015 over 2014

Given

GDP 2014 = 2 trillion $

GDP 2015 = 3 trillion $

GDP Growth = (3-2)/2 = 50 % growth in GDP over previous year

GDP = Total market value of the goods and services produced within a country within a given time period.

Inflation Rate :

Inflation Rate = [CPI(2015) - CPI (2014)]/CPI(2014)

Suppose the Consumer Price Index of a basket of goods is given by :

CPI 2014 = 50 $

CPI 2015 = 100 $

Inflation Rate : 100-50/50 = 100 % inflation , i.e the prices have doubled of a basket of goods .

High inflation rate is undesirable.

CPI = Consumer Price Index of a specific basket of goods within a country.

Exchange Rate :

C(A)/C(B)

where currency C(A) is the value of our currency(south african) in terms of USD $

and C(B) is one USD $

suppose the value of 1 USD $ = 14.65 South african rand

Therefore Exchange Rate : 1/14.65 = 0.068 rand/USD $

A low exchange rate signifies low strength of the currency .

BoP : Net Current Account + Net Capital Account + Net Financial Account =

Net Current Account = (Net Exports - Net Imports)

Greater the amount of exports and lesser the imports better the economy .

Net Capital Account = ( Capital Inflows - Capital Outflows)

Net inflows represents foreign ownership within the country i.e money is flowing in the country .

Net Outflows represents country ownership in foreign i.e money is flowing out of the country .

Net Financial Account = (Borrowings - Investments)

Borrowings : Loan taken from other countries , economies , IMF , World Bank etc

Investments : Investments in the form of FDI .

GDP per capita :